OpenStudy (bohotness):

Help me please k12 7 grade really help

OpenStudy (bohotness):

@undeadknight26 @ganeshie8 @Preetha @thomaster @TheAsker2002 @Deadpool. @demonchild99

OpenStudy (bohotness):

@confluxepic @Compassionate @curlette123

OpenStudy (bohotness):

Zach solved this problem. Maria had a spool of ribbon 68 inches long. She cut the ribbon into 8 equal-length pieces. How long was each piece of ribbon? Zach said the answer was 0.85 inches. Is Zach's answer reasonable? A. Yes B. No, because 64 ÷ 8 = 8 and 72 ÷ 8 = 9, so the answer should be between 8 and 9. C. No, because 70 + 8 = 78, and Maria has less than 70 inches of ribbon, so the answer should be a bit less than 78. D. No, because 70 • 8 = 560 and Maria has less than 70 inches of ribbon, so the answer should be a bit less than 560.

OpenStudy (bohotness):

@jabez177

jabez177 (jabez177):

sorry never did this or i forgot

OpenStudy (bohotness):

ok i am struck on unit4 lesson 8

jabez177 (jabez177):

what subject

OpenStudy (bohotness):

math

jabez177 (jabez177):

ok let me check

OpenStudy (bohotness):

ok

OpenStudy (bohotness):

jues glitcjh my problem changed to this

OpenStudy (bohotness):

Convert the fraction to a decimal. A. 0.16 B. 0.425 C. 1.6 D. 4.25

OpenStudy (bohotness):

i picked d

jabez177 (jabez177):

yeah sorry cant help gtg

OpenStudy (bohotness):

ok

OpenStudy (anonymous):

what fraction? there is not a fraction here!!!

OpenStudy (bohotness):

Convert the fraction to a decimal. A. B. C. D.

OpenStudy (bohotness):

thats number two

OpenStudy (bohotness):

hold on i post 1s faction

OpenStudy (bohotness):

that one

OpenStudy (bohotness):

this goes to nubmer 2

OpenStudy (bohotness):

Convert the fraction to a decimal. A. B. C. D.

OpenStudy (anonymous):

oh well, I can't help. @Hero Help!!!

OpenStudy (bohotness):

Convert the fraction to a decimal. A. B. C. D.

OpenStudy (bohotness):

Which mixed number is equivalent to this decimal? 9.003 A. B. C. D.